OpenStudy (anonymous):

Sandra plans to fill a corner of her yard with decorative stones. Each row has seven more stones than the prior row. If there were five stones in the first row, how many stones were used to fill all 36 rows? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

okay, back to this problem. What this is a sum problem. Now i am sure you are familiar with the formula: Sum=(n/2)(first +last)=(n/2)[2(5)+(n-1)*d

OpenStudy (anonymous):

now, our first row had 5 stones

OpenStudy (anonymous):

now our common difference is of course 7, because we add 7 to every row after the first. So : d=7

OpenStudy (anonymous):

no in this case n=36, because we want 36 rows

OpenStudy (anonymous):

now we plug our information into the formula, and we get: Sum=36/2[10+(36-1)*7 now all we do is evluate that: sum=4590

OpenStudy (anonymous):

that is our answer , 4590 stones